For the synthesis of butylethanoate 11x10^-3 mol of the alcohol and 13x10^-3 mol of carboxylic acid were used. Assuming that you obtained 6.6x10^-3 mol of butylethanoate, calculate the percentage yield.
1
Expert's answer
2020-03-25T12:47:47-0400




find the amount of substance theoretical product


count on alcohol, it reacted less

 11x10-3:1=х:1,

х=11x10-3.

calculate the percentage yield.


"v=\frac{vp}{vt}=\frac{6.6\times10^-3}{11\times10^-3}=0.6"

"0.6\times100=60" , 60%